# Upstream history quantification and scale-decomposed energy analysis for weak-to-strong adverse-pressure-gradient turbulent boundary layers

### v1

@@ Abstract
The present study delineates the effects of pressure gradient history and local disequilibration on the small and large-scale energy in turbulent boundary layers (TBLs) imposed with a broad range of adverse-pressure-gradients (APG). This is made possible by analyzing four published high-fidelity APG TBL databases, which span weak to strong APGs and cover dynamic conditions ranging from near-equilibrium to strong disequilibrium. These databases enable the development of a methodology to understand the effects of PG history and local disequilibration, the latter defined here as the local streamwise rate of change of the pressure force contribution in the force balance. The influence of PG history on TBL statistics is quantified by the accumulated PG parameter ($\overline{\beta}$), proposed previously by Vinuesa et al. (2017) to study integral quanitites, which is compared here between cases at matched local PG strength ($\beta$), Reynolds number ($Re$) and ${\rm d}{\beta}/{{\rm d}{Re}}$ at nominally similar orders of magnitude. While the effects of local disequilibration (${\rm d}{\beta}/{{\rm d}{Re}}$) are investigated by considering TBL cases at matched $\beta$, $Re$, and fairly matched $\overline{\beta}$. This enables analysis of accumulated PG history and local disequilibration effects separately, where applicable, to highlight qualitative differences in statistical trends. It is found that $\overline{\beta}$ cannot unambiguously capture history effects when ${\rm d}{\beta}/{{\rm d}{Re}}$ levels are significantly high, as it does not account for the delayed response of the mean flow and turbulence, nor the attenuation of the pressure gradient effect with distance. In two comparisons of APG TBLs under strong non-equilibrium, the values of $\overline{\beta}$ and ${\rm d}{\beta}/{\rm d}{Re}$ expressed using Zagarola-Smits scaling were found to be consistent with the trends in mean velocity defect and Reynolds stresses noted previously for weak APG TBLs. While an increase in $\overline{\beta}$ is associated with energisation of both the small and large scales in the outer regions of APG TBLs, it affects only the large scales in the near-wall region. This confirms the ability of near-wall small scales to rapidly adjust to changes in PG strength. By attempting to provide a structured parametric methodology to isolate effects of PG history and local disequilibration, this study reports the influence of these effects on turbulent flow statistics across the widest APG strengths documented in the literature.
@@ 1 Introduction
Adverse-pressure-gradient (APG) turbulent boundary layers (TBLs) are encountered in various settings such as the flow over wings, automobiles and in turbomachinery applications. Hence, understanding the effects of pressure gradients on TBLs is crucial for developing flow control and drag reduction strategies for improving efficiencies of these engineering applications. Equilibrium TBLs represent the simplest but rarest forms of PG TBLs, characterized by a constant force balance as the flow develops (Rotta, 1950; Clauser, 1956; Townsend, 1956). Here, equilibrium refers to the unique state of the TBL where the flow properties remain dynamically similar despite the flow developing/evolving in the streamwise direction (Maciel et al., 2006; Devenport and Lowe, 2022). Achieving equilibrium in TBLs is, however, challenging due to its complex multi-scale behavior, with each layer of the flow governed by characteristic length/time scales and force balances. The concept of equilibrium is inherently related with self-similarity; a flow is said to be self-similar if its statistical properties are solely influenced by local flow variables (Kitsios et al., 2017). However, almost all the APG TBLs in practise exist in a non-equilibrium state, where turbulence statistics exhibit a multi-parameter dependence on local PG strength, flow Reynolds numbers, upstream PG/history as well as local disequilibration effects (Bobke et al., 2017; Gungor et al., 2024a).
Historically, the first steps in understanding PG effects on TBL statistics were focused on quantification of the local PG strength, such as by the Clauser parameter $\beta_{C}$ (Clauser, 1956), involving estimation of the ratio of pressure force to a reference force corresponding to the TBL. Per definition, $\beta_{C}=$
@@
$(\delta^{*}/\tau_{w})(dP/dx)$, where $\delta^{*}$ is the TBL displacement thickness, $\tau_{w}$ is the mean wall shear stress and $(dP/dx)$ is the mean pressure-gradient over the streamwise distance x. Specifically, $\beta_{C}$ quantifies the relative strength of the streamwise pressure force to the wall shear force, providing a global measure of the force balance across the boundary layer. Past studies have noted the variation in pressure gradient strength to lead to notable deviations in turbulent flow statistics from classical scaling theories, such as the log-law scaling of the mean streamwise velocity (Coles, 1956; Marusic et al., 2013; Knopp et al., 2021). Additionally, APG TBLs are characterized by greater mean velocity defects and increased intensities of Reynolds stress components, particularly in their outer regions (Skåre and Krogstad, 1994; Harun et al., 2013). Notably, all the Reynolds stresses in APG TBLs display a distinctive outer peak at low friction Reynolds numbers, which, in ZPG TBLs are expected to appear only at very high $Re_{\tau}$ (Willert et al., 2017; Deshpande and Vinuesa, 2024). Here the friction Reynolds number is defined as $Re_{\tau}=U_{\tau}\delta/\nu$, where $U_{\tau}$ is the mean friction velocity, $\delta$ is the mean boundary layer thickness, and $\nu$ is the kinematic viscosity. The association of these intensified Reynolds stresses with their corresponding Reynolds transport equations, for varying APG strengths, has been discussed previously by Gungor et al. (2022) and Deshpande and Vinuesa (2024). On the other hand, the effects of $Re_{\tau}$ on the outer-region energy of APG TBLs have been investigated by Monty et al. (2011), Pozuelo et al. (2022) and Deshpande et al. (2023). Specifically, Deshpande et al. (2023) employed an experimental setup that minimized the effects of history and disequilibration by controlling the airflow from the wind tunnel ceiling. They also analyzed the small- and large-scale contributions to the Reynolds stresses in weak-to-moderately strong APGs, which unraveled the decrease in their small-scale energy for an increase in $Re_{\tau}$ (Deshpande et al., 2023). Notably, this $Re_{\tau}$-trend is in contrast with those observed for increasing APG magnitudes and upstream history effects, which past studies have held responsible for increasing both the small- and large-scale energies (Bobke et al., 2017; Monty et al., 2011; Tanarro et al., 2020). However, it still remains unknown if PG history influences the small and large-scale energies in the same way for strong APG cases, which this study will investigate.
@@
To summarize this section, despite extensive research on the effects of APG in the past (as briefly introduced above), no study has isolated the impact of upstream PG history as well as local disequilibration across weak-to-strong APG strengths. For instance, although Bobke et al. (2017) isolated the effects of history by focusing on matched $\beta_{C}$ and $Re_{\tau}$ cases, the analysis was limited to low-$\beta_{C}$ TBLs. On the other hand, Gungor et al. (2024b) investigated the effects of history on non-equilibrium APG TBLs with high $\beta$, but without isolating the Reynolds number and local disequilibration effects. What is thus lacking is a comprehensive analysis that quantifies the effects of upstream history and local disequilibration over a wide range of $\beta$ parameters based on a choice of consistent scaling parameters. This will be attempted here by drawing inspiration from the accumulated $\beta$-averaging method of Vinuesa et al. (2017), which will be imposed on the alternative PG parameters proposed by studies discussed above. This forms the primary focus of the present study, which aims to propose a structured methodology to understand and quantify the effects of upstream PG history and local disequilibration. Inspired by the work of Vinuesa et al. (2017) and Gungor et al. (2024b), we test using $\overline{\beta}_{ZS}$, defined as an average of $\beta_{ZS}$ over $Re_{ZS}$ or $Re_{\theta}$, to quantify accumulated outer-region pressure gradients for strong APG TBLs at matched $\beta_{ZS}$ and $Re_{ZS}$ or $Re_{\theta}$. Similarly, we test $\overline{\beta}_{C,\theta}$, defined as an average of $\beta_{C}$ over $Re_{\theta}$, for weak APG TBLs at matched $\beta_{C}$ and $Re_{\theta}$. It must be noted that these accumulated PG parameters, by construction, do not fully capture the flow history as they do not account for delayed turbulence response or the attenuation of pressure gradient effects with streamwise distance. Consequently, particular attention will be given to assessing the significance of this limitation. Furthermore, the primary reason behind using $\beta_{ZS}$ instead of $\beta_{C}$ lies in the fact that $\beta_{C}$ does not capture the force balance for strong APG cases, whereas $\beta_{ZS}$ remains valid across all APG cases. In the present work, we employ five published, high-fidelity ZPG and APG datasets (Bobke et al., 2017; Pozuelo et al., 2022; Gungor et al., 2016, 2022) covering a wide range of $\beta$ and Reynolds numbers. Two of these four datasets correspond to near-equilibrium TBLs with weak to moderate APGs, while the other two are APG TBLs in strong non-equilibrium. Velocity variances and mean velocity contributions will be analyzed to assess the applicability of the history and local disequilibration parameters across broad APG strengths and equilibrium states. In addition to offering several cases with matched $\beta$ and $Re$ (for appropriate choice of scaling), the datasets include a unique case where $\beta$ and $Re$ are matched across three TBLs, allowing us to closely study the effects of local disequilibration and accumulated PG. Once established, the varying upstream history and disequilibration cases are used to understand effects on small- and large-scale velocity variance, achieved via scale-based energy decomposition.
@@
Figure 1: $\beta$ vs $Re$ plots illustrating the Exactly-Matched and Quasi-Matched cases across the four data sets. Exactly-Matched cases for (a) $\beta_{C}$ vs $Re_{\theta}$ ($\beta_{C}$ vs $Re_{\tau}$ in the inset), (b) $\beta_{ZS}$ vs $Re_{\theta}$ and (c) $\beta_{ZS}$ vs $Re_{ZS}$ are shown using symbols: $\circ$- Case 1; $\square$- Case 2; $\diamond$- Case 3; $\bigtriangleup$- Case 4 and $\bigtriangledown$- Case 5. Line plots correspond to datasets mentioned in Table 1. The white-filled symbols indicate the Exactly-Matched cases while the symbols filled in black represent Quasi-matched cases, which were selected due to the unavailability of spectral data at the location of the white-filled symbols. Further details about these cases can be found in Table 2 and in §$\S$2.2.
@@
Case# Dataset Line Style Exactly-Matched $\beta$, $Re$ Quasi-Matched $\beta$, $Re$
@@
Case 5 DNS22 3250 0.16 0.25 -3.18e-05 2.10 3760.4 0.15 0.23 -2.28e-05 2.15
DNS16 3250 0.16 0.31 -6.96e-05 2.50 3467.1 0.15 0.30 -6.63e-05 2.56
Table 2: Summary of the various cases analyzed in this study, with Case# corresponding to $\beta$ and $Re$ marked in Figure 1. The "Exactly-Matched" subgroup indicates the parametric values for comparison of mean velocities and variances in subsequent analysis, while "Quasi-Matched" subgroup contains the parametric values considered for comparing the scale-decomposed variances, due to the unavailability of spectral energy data at "Exactly-matched" cases. Only relevant parameters for each case are included. Different line styles have been considered for Cases 3 to 5 to enhance comprehension in Figures 4 and 5.
@@ 2.2 Case Selection and Analysis Methodology
Here, we draw inspiration from Bobke et al. (2017) and use the four APG TBL databases to identify matched $\beta$ and Re cases to limit the differences between them solely to upstream PG history and local disequilibrating effects. The selected cases (illustrated in Figure 1) are based on three differently scaled Reynolds numbers ($Re_{\tau}$, $Re_{\theta}$, and $Re_{ZS}$) and two pressure-gradient parameters ($\beta_{C}$ and $\beta_{ZS}$) to assess their ability in truly isolating the aforementioned effects (Bobke et al., 2017; Maciel et al., 2006; Gungor et al., 2024b, a). Table 2 documents the values of relevant parameters for the "Exactly-Matched" and "Quasi-Matched" cases considered for the present analysis, with the differences between them obvious from the nomenclature. While the former are considered for comparing mean velocities and variances in the subsequent analysis, the latter subgroup is considered for comparing the scale-decomposed variances due to the unavailability of spectral energy data for the "Exactly-matched" cases (which are typically only saved at certain streamwise locations of the domain).
@@
Case 3 compares relatively downstream locations of the numerical domains associated with b1.4 and DNS22, while Case 4 compares the two high-$\beta$ non-equilibrium datasets in their relatively upstream regions, where the local pressure force impact, quantified by $\beta_{ZS}$, is high. Case 5 allows us to compare two non-equilibrium boundary layers at matched $\beta_{ZS}$ and $Re_{ZS}$ values, hence utilizing the Zagarola-Smits scaling for both pressure-gradient and Reynolds number. Here, $Re_{ZS}$ and $\beta_{ZS}$ are obtained based on a momentum equation analysis demonstrated previously in Maciel et al. (2018), where the order of magnitude of the Reynolds shear stress gradient term in the outer region is kept constant in the momentum equation. The choice of the $U_{ZS}$ velocity scale results in $Re_{ZS}$ and $\beta_{ZS}$ as non-dimensional parameters that accurately follow the force ratios in the outer region.
It must be noted that although here we are using $Re_{\theta}$ as a suitable parameter for TBLs with large-defects, it does not have a rigorous theoretical formulation, unlike $Re_{ZS}$. The similarities and differences observed (if any) between matched cases using these two different Reynolds number definitions will be discussed in §$\S$3. Note that for Cases 3 and 4, where the cases are matched based on $Re_{\theta}$, the definition for the accumulated $\beta_{ZS}$ follows equation (2), while Case 5 being a matched case of $Re_{ZS}$ is defined by equation (3) following:
@@
Averaging in this way aims at capturing the nature of upstream development of the APG TBLs over the relevant Reynolds number range, thus quantifying their accumulated history.
Availability of spectral energy data at the "Quasi-Matched" cases plays a crucial role in this study, enabling analysis of small- and large-scale energy contributions as a function of upstream PG history. Although the $\beta$ and $Re$ values do not match exactly for these comparisons, the slight mismatch does not influence the statistical trends inferred in subsequent sections. We utilize the cutoff based on the viscous-scaled spanwise wavelength ($\lambda^{+}_{z,c}=300$) proposed by Deshpande et al. (2023) to decompose the total spectral energy at each wall-normal location into small ($\overline{u^{2}_{SS}}$; $\lambda^{+}_{z,c}$leq 300) and large scale contributions ($\overline{u^{2}_{LS}}$; $\lambda^{+}_{z,c}$ $>$ 300), such that $\overline{u^{2}}$($y^{+}$) = $\overline{u^{2}_{SS}}$($y^{+}$) + $\overline{u^{2}_{LS}}$($y^{+}$), and $\overline{v^{2}}$($y^{+}$) = $\overline{v^{2}_{SS}}$($y^{+}$) + $\overline{v^{2}_{LS}}$($y^{+}$). Deshpande et al. (2023) arrived at this cutoff value iteratively, based on the well-accepted argument of viscous-scaled small-scale energy ($\overline{u^{2}_{SS}}^{+}$) exhibiting $Re_{\tau}$-invariance for $\lambda_{z,c}^{+}\leq 300$ (for the case of canonical flows; refer to figure 12 of Deshpande et al., 2023). Since this spanwise wavelength cutoff, $\lambda^{+}_{z,c}$ is constructed based on a viscous-scaled parameter that breaks down for strong APG TBLs, it becomes ineffective in higher-defect cases. This limitation is overcome by utilizing an outer-scaled spanwise wavelength ($\lambda_{z,c}/\delta$), the cutoff value for which is estimated iteratively such that the small- and large-scale velocity variance profiles nominally match those computed based on the established value of $\lambda^{+}_{z,c}$ = 300, for the low-defect Cases 1 and 2. The cutoff value of $\lambda_{z,c}/\delta$ = 0.5 shows reasonable agreement between the scale-decomposed variances, as demonstrated in the plots provided in Appendix A.

### v2

@@ Abstract
The present study delineates the effects of pressure gradient history and local disequilibration on the small and large-scale energy in turbulent boundary layers (TBLs) imposed with a broad range of adverse-pressure-gradients (APG). This is made possible by analyzing four published high-fidelity APG TBL databases, which span weak to strong APGs and cover dynamic conditions ranging from near-equilibrium to strong disequilibrium. These databases enable the development of a methodology to understand the effects of PG history and local disequilibration, the latter defined here as the local streamwise rate of change of the pressure force contribution in the force balance. The influence of PG history on TBL statistics is quantified by the accumulated PG parameter ($\overline{\beta}$), proposed previously by Vinuesa et al. (2017) to study integral quanitites, which is compared here between cases at matched local PG strength ($\beta$), Reynolds number ($Re$) and ${\rm d}{\beta}/{{\rm d}{Re}}$ at nominally similar orders of magnitude. Here, $\beta$ denotes a general umbrella term used for pressure gradient parameters that is estimated using different scaling parameters in this study. While the effects of local disequilibration (${\rm d}{\beta}/{{\rm d}{Re}}$) are investigated by considering TBL cases at matched $\beta$, $Re$, and fairly matched $\overline{\beta}$. This enables analysis of accumulated PG history and local disequilibration effects separately, where applicable, to highlight qualitative differences in statistical trends. It is found that $\overline{\beta}$ cannot unambiguously capture history effects when ${\rm d}{\beta}/{{\rm d}{Re}}$ levels are significantly high, as it does not account for the delayed response of the mean flow and turbulence, nor the attenuation of the pressure gradient effect with distance. In two comparisons of APG TBLs under strong non-equilibrium, the values of $\overline{\beta}$ and ${\rm d}{\beta}/{\rm d}{Re}$ expressed using Zagarola-Smits scaling were found to be consistent with the trends in mean velocity defect and Reynolds stresses noted previously for weak APG TBLs. While an increase in $\overline{\beta}$ is associated with energisation of both the small and large scales in the outer regions of APG TBLs, it affects only the large scales in the near-wall region. This confirms the ability of near-wall small scales to rapidly adjust to changes in PG strength. By attempting to provide a structured parametric methodology to isolate effects of PG history and local disequilibration, this study reports the influence of these effects on turbulent flow statistics across the widest APG strengths documented in the literature.
@@ 1 Introduction
Adverse-pressure-gradient (APG) turbulent boundary layers (TBLs) are encountered in various settings such as the flow over wings, automobiles and in turbomachinery applications. Hence, understanding the effects of pressure gradients on TBLs is crucial for developing flow control and drag reduction strategies for improving efficiencies of these engineering applications. Equilibrium TBLs represent the simplest but rarest forms of PG TBLs, characterized by a constant force balance as the flow develops (Rotta, 1950; Clauser, 1956; Townsend, 1956). Here, equilibrium refers to the unique state of the TBL where the flow properties remain dynamically similar despite the flow developing/evolving in the streamwise direction (Maciel et al., 2006; Devenport and Lowe, 2022). Achieving equilibrium in TBLs is, however, challenging due to its complex multi-scale behavior, with each layer of the flow governed by characteristic length/time scales and force balances. The concept of equilibrium is inherently related with self-similarity; a flow is said to be self-similar if its statistical properties are solely influenced by local flow variables (Kitsios et al., 2017). However, almost all the APG TBLs in practise exist in a non-equilibrium state, where turbulence statistics exhibit a multi-parameter dependence on local PG strength, flow Reynolds numbers, upstream PG/history as well as local disequilibration effects (Bobke et al., 2017; Gungor et al., 2024a). Historically, the first steps in understanding PG effects on TBL statistics were focused on quantification of the local PG strength, such as by the Clauser parameter $\beta_{C}$ (Clauser, 1956), involving estimation of the ratio of pressure force to a reference force corresponding to the TBL. Per definition, $\beta_{C}=$ $(\delta^{*}/\tau_{w})(dP/dx)$, where $\delta^{*}$ is the TBL displacement thickness, $\tau_{w}$ is the mean wall shear stress and $(dP/dx)$ is the mean pressure-gradient over the streamwise distance x.
Nomenclature Symbols $C_{f}$ Skin friction coefficient $H_{12}$ Shape factor, $\delta^{*}/\theta$ $l_{\tau}$ Friction-viscous length scale, $\nu/U_{\tau}$ $Re_{\tau}$ Friction Reynolds number, $U_{\tau}\delta/\nu$ $Re_{\theta}$ Reynolds number based on $\theta$, $U_{e}\theta/\nu$ $Re_{ZS}$ Zagarola-Smits Reynolds number $U$ Mean streamwise velocity $U_{e}$ Boundary layer edge velocity $U_{\tau}$ Friction velocity $U_{ZS}$ Zagarola-Smits velocity scale $V$ Mean wall-normal velocity $W$ Mean spanwise velocity $u,v,w$ Velocity fluctuations in $x,y,z$ directions $\overline{u^{2}}$, $\overline{v^{2}}$, $w^{2}$ Reynolds normal stresses $u^{2}_{LS}$ Large-scale contribution to $\overline{u^{2}}$ $u^{2}_{SS}$ Small-scale contribution to $\overline{u^{2}}$ $v^{2}_{LS}$ Large-scale contribution to $\overline{v^{2}}$ $v^{2}_{SS}$ Small-scale contribution to $\overline{v^{2}}$ $x$ Streamwise coordinate $y$ Wall-normal coordinate $z$ Spanwise coordinate Greek Symbols $\beta_{C}$ Clauser pressure-gradient parameter $\beta_{ZS}$ Zagarola-Smits pressure-gradient parameter ${d\beta}/{dRe}$ Local disequilibration parameter $\overline{\beta}$ Accumulated pressure-gradient parameter $\delta$ Boundary layer thickness $\delta^{*}$ Displacement thickness $\lambda_{z,c}$ Spanwise wavelength cutoff for scale decomposition $\nu$ Kinematic viscosity $\phi_{uu}$ Power spectral density of $u$ $\phi_{vv}$ Power spectral density of $v$ $\rho$ Fluid density $\tau_{w}$ Wall shear stress $\theta$ Momentum thickness Subscripts and Superscripts $()_{0}$ Value under reference or neutral condition $\overline{()^{2}}$ Variances $\overline{()}$ Time or ensemble average Abbreviations APG Adverse pressure gradient DNS Direct numerical simulation LES Large-eddy simulation TBL Turbulent boundary layer ZPG Zero pressure gradient
@@
Specifically, $\beta_{C}$ quantifies the relative strength of the streamwise pressure force to the wall shear force, providing a global measure of the force balance across the boundary layer. Past studies have noted the variation in pressure gradient strength to lead to notable deviations in turbulent flow statistics from classical scaling theories, such as the log-law scaling of the mean streamwise velocity (Coles, 1956; Marusic et al., 2013; Knopp et al., 2021). Additionally, APG TBLs are characterized by greater mean velocity defects and increased intensities of Reynolds stress components, particularly in their outer regions (Skåre and Krogstad, 1994; Harun et al., 2013). Notably, all the Reynolds stresses in APG TBLs display a distinctive outer peak at low friction Reynolds numbers, which, in ZPG TBLs are expected to appear only at very high $Re_{\tau}$ (Willert et al., 2017; Deshpande and Vinuesa, 2024). Here the friction Reynolds number is defined as $Re_{\tau}=U_{\tau}\delta/\nu$, where $U_{\tau}$ is the mean friction velocity, $\delta$ is the mean boundary layer thickness, and $\nu$ is the kinematic viscosity. The association of these intensified Reynolds stresses with their corresponding Reynolds transport equations, for varying APG strengths, has been discussed previously by Gungor et al. (2022) and Deshpande and Vinuesa (2024). On the other hand, the effects of $Re_{\tau}$ on the outer-region energy of APG TBLs have been investigated by Monty et al. (2011), Pozuelo et al. (2022) and Deshpande et al. (2023). Specifically, Deshpande et al. (2023) employed an experimental setup that minimized the effects of history and disequilibration by controlling the airflow from the wind tunnel ceiling. They also analyzed the small- and large-scale contributions to the Reynolds stresses in weak-to-moderately strong APGs, which unraveled the decrease in their small-scale energy for an increase in $Re_{\tau}$ (Deshpande et al., 2023). Notably, this $Re_{\tau}$-trend is in contrast with those observed for increasing APG magnitudes and upstream history effects, which past studies have held responsible for increasing both the small- and large-scale energies (Bobke et al., 2017; Monty et al., 2011; Tanarro et al., 2020). However, it still remains unknown if PG history influences the small and large-scale energies in the same way for strong APG cases, which this study will investigate.
@@
To summarize this section, despite extensive research on the effects of APG in the past (as briefly introduced above), no study has isolated the impact of upstream PG history as well as local disequilibration across weak-to-strong APG strengths. For instance, although Bobke et al. (2017) isolated the effects of history by focusing on matched $\beta_{C}$ and $Re_{\tau}$ cases, the analysis was limited to low-$\beta_{C}$ TBLs. On the other hand, Gungor et al. (2024b) investigated the effects of history on non-equilibrium APG TBLs with high $\beta$, but without isolating the Reynolds number and local disequilibration effects. What is thus lacking is a comprehensive analysis that quantifies the effects of upstream history and local disequilibration over a wide range of $\beta$ parameters based on a choice of consistent scaling parameters. This will be attempted here by drawing inspiration from the accumulated $\beta$-averaging method of Vinuesa et al. (2017), which will be imposed on the alternative PG parameters proposed by studies discussed above. This forms the primary focus of the present study, which aims to propose a structured methodology to understand and quantify the effects of upstream PG history and local disequilibration. Inspired by the work of Vinuesa et al. (2017) and Gungor et al. (2024b), we test using $\overline{\beta}_{ZS}$, defined as an average of $\beta_{ZS}$ over $Re_{ZS}$ or $Re_{\theta}$, to quantify accumulated outer-region pressure gradients for strong APG TBLs at matched $\beta_{ZS}$ and $Re_{ZS}$ or $Re_{\theta}$. Similarly, we test $\overline{\beta}_{C,\theta}$, defined as an average of $\beta_{C}$ over $Re_{\theta}$, for weak APG TBLs at matched $\beta_{C}$ and $Re_{\theta}$. It must be noted that these accumulated PG parameters, by construction, do not fully capture the flow history as they do not account for delayed turbulence response or the attenuation of pressure gradient effects with streamwise distance. Consequently, particular attention will be given to assessing the significance of this limitation. Furthermore, the primary reason behind using $\beta_{ZS}$ instead of $\beta_{C}$ lies in the fact that $\beta_{C}$ does not capture the force balance for strong APG cases, whereas $\beta_{ZS}$ remains valid across all APG cases.
In the present work, we employ five published, high-fidelity ZPG and APG datasets (Bobke et al., 2017; Pozuelo et al., 2022; Gungor et al., 2016, 2022) covering a wide range of $\beta$ and Reynolds numbers. Two of these four datasets correspond to near-equilibrium TBLs with weak to moderate APGs, while the other two are APG TBLs in strong non-equilibrium. Velocity variances and mean velocity contributions will be analyzed to assess the applicability of the history and local disequilibration parameters across broad APG strengths and equilibrium states. In addition to offering several cases with matched $\beta$ and $Re$ (for appropriate choice of scaling), the datasets include a unique case where $\beta$ and $Re$ are matched across three TBLs, allowing us to closely study the effects of local disequilibration and accumulated PG. Once established, the varying upstream history and disequilibration cases are used to understand effects on small- and large-scale velocity variance, achieved via scale-based energy decomposition. Note that the symbol $\beta$ is used as a generic or umbrella notation for local pressure-gradient (PG) measures, with $\beta_{C}$ and $\beta_{ZS}$ representing specific scaled forms under this general definition.
@@
Figure 1: $\beta$ vs ReRe plots illustrating the Exactly-Matched and Quasi-Matched cases across the four data sets. Exactly-Matched cases for (a) $\beta_{C}$ vs $Re_{\theta}$ ($\beta_{C}$ vs $Re_{\tau}$ in the inset), (b) $\beta_{ZS}$ vs $Re_{\theta}$ and (c) $\beta_{ZS}$ vs $Re_{ZS}$ are shown using symbols: $\circ$- Case 1; $\square$- Case 2; $\diamond$- Case 3; $\bigtriangleup$- Case 4 and $\bigtriangledown$- Case 5. Filled cyan circles indicate Case DISEQ. Line plots correspond to datasets mentioned in Table 1. The white-filled symbols indicate the Exactly-Matched cases while the symbols filled in black represent Quasi-matched cases, which were selected due to the unavailability of spectral data at the location of the white-filled symbols. Further details about these cases can be found in Table 2 and in $\S$$2.$
@@
Case# Dataset Line Style Exactly-Matched $\beta$, ReRe Quasi-Matched $\beta$, ReRe
@@
Case 5 DNS22 3250 0.16 0.25 -3.2e-05 2.10 3760.4 0.15 0.23 -2.3e-05 2.15
DNS16 3250 0.16 0.31 -7e-05 2.50 3467.1 0.15 0.30 -6.6e-05 2.56
Case DISEQ : Not Matched (included for comparative analysis)
Dataset Line Style $\mathbf{Re_{\theta}}$ $\mathbf{\beta_{C}}$ $\mathbf{\overline{\beta}_{C,\theta}}$ $\mathbf{d\beta_{C}/dRe_{\theta}}$ $\mathbf{H_{12}}$
Case DISEQ DNS22 2170 1.72 0.89 2.9e-03 1.50
b1.4 3098 1.59 0.89 2.2e-04 1.61
Table 2: Summary of the various cases analyzed in this study, with Case# corresponding to $\beta$ and ReRe marked in Figure 1. The "Exactly-Matched" subgroup indicates the parametric values for comparison of mean velocities and variances in subsequent analysis, while "Quasi-Matched" subgroup contains the parametric values considered for comparing the scale-decomposed variances, due to the unavailability of spectral energy data at "Exactly-matched" cases. Only relevant parameters for each case are included. Different line styles have been considered for Cases 3 to 5 to enhance comprehension in Figures 5 and 6. Note that Case DISEQ is not a matched case and is included to study the effects of local disequilibration on turbulent scales.
@@ 2.2 Case Selection and Analysis Methodology
Here, we draw inspiration from Bobke et al. (2017) and use the four APG TBL databases to identify matched $\beta$ and Re cases to limit the differences between them solely to upstream PG history and local disequilibrating effects. The selected cases (illustrated in Figure 1) are based on three differently scaled Reynolds numbers ($Re_{\tau}$, $Re_{\theta}$, and $Re_{ZS}$) and two pressure-gradient parameters ($\beta_{C}$ and $\beta_{ZS}$) to assess their ability in truly isolating the aforementioned effects (Bobke et al., 2017; Maciel et al., 2006; Gungor et al., 2024b, a). Table 2 documents the values of relevant parameters for the "Exactly-Matched" and "Quasi-Matched" cases considered for the present analysis, with the differences between them obvious from the nomenclature. While the former are considered for comparing mean velocities and variances in the subsequent analysis, the latter subgroup is considered for comparing the energy spectra and scale-decomposed variances due to the unavailability of spectral energy data for the "Exactly-matched" cases (which are typically only saved at certain streamwise locations of the domain).
@@
Case 3 compares relatively downstream locations of the numerical domains associated with b1.4 and DNS22, while Case 4 compares the two high-$\beta$ non-equilibrium datasets in their relatively upstream regions, where the local pressure force impact, quantified by $\beta_{ZS}$, is high. Case 5 allows us to compare two non-equilibrium boundary layers at matched $\beta_{ZS}$ and $Re_{ZS}$ values, hence utilizing the Zagarola-Smits scaling for both pressure-gradient and Reynolds number. For reference, the $Re_{\theta}$ values corresponding to the Exactly-Matched column in Case 5 are 4481 for DNS22 and 3175 for DNS16, while the Quasi-Matched column corresponds to 4737 and 3229, respectively.
Case DISEQ represents a special scenario with fairly matched $\beta_{C}$ values, identical $\overline{\beta}_{C,\theta}$ values, and local disequilibration values that differ by an order of magnitude. While the $Re_{\theta}$ values differ (2200 for DNS22 and 3100 for b1.4), trends in the small-scale near-wall region can be attributed to local disequilibration due to the region’s sensitivity to PG changes. Case DISEQ represents a location that is not matched in $\beta$ and $Re$, but enables a more focused examination of local disequilibration effects on turbulent scales. An in-depth discussion on the power-spectral density in Case DISEQ is provided in Section 3.1.
Note that $Re_{ZS}$ and $\beta_{ZS}$ are obtained based on a momentum equation analysis demonstrated previously in Maciel et al. (2018), where the order of magnitude of the Reynolds shear stress gradient term in the outer region is kept constant in the momentum equation. The choice of the $U_{ZS}$ velocity scale results in $Re_{ZS}$ and $\beta_{ZS}$ as non-dimensional parameters that accurately follow the force ratios in the outer region.
It must be noted that although here we are using $Re_{\theta}$ as a suitable parameter for TBLs with large-defects, it does not have a rigorous theoretical formulation, unlike $Re_{ZS}$. The similarities and differences observed (if any) between matched cases using these two different Reynolds number definitions will be discussed in $\S$3. Note that for Cases 3 and 4, where the cases are matched based on $Re_{\theta}$, the definition for the accumulated $\beta_{ZS}$ follows equation (2), while Case 5 being a matched case of $Re_{ZS}$ is defined by equation (3) following:
@@
Averaging in this way aims at capturing the nature of upstream development of the APG TBLs over the relevant Reynolds number range, thus quantifying their accumulated history. For the b1.0 and b1.4 datasets, the lower limit $Re_{\theta,0}$ is chosen at a slightly downstream location from the start of their simulation domains ($Re_{\theta}$ $\sim$ 320320), to ensure that the boundary layer is fully turbulent and under ZPG condition. On the other hand, this lower limit is at the start of the DNS22 and DNS16 simulation domains, where the boundary layer is already fully turbulent.